# Towards Assessing Medical Ethics from Knowledge to Practice

### v1

@@ Appendix A Data Source
For open-ended questions, case materials were systematically collected from the Case and Commentary, AMA Journal of Ethics. (American Medical Association, 1999–2025), covering all publications from January 1, 1999, to June 30, 2025.
Appendix B Sample of Curated MCQA
MCQAs in PrinciplismQA is synthesized from content excerpts highlighting the relevant ethical concept(s). The structure of a curated question focused on clinical or professional application, four answer options (with one correct answer), and an explanation that justifies the correct choice and addresses why the other options are incorrect. Figure 4 provides an example MCQA curation task.
Appendix C Intraclass Correlation Coefficient (ICC) Calculation Formula

### v2

@@ Appendix C Data Source
For open-ended questions, case materials were systematically collected from the Case and Commentary, AMA Journal of Ethics. American Medical Association (1999), covering all publications from January 1, 1999, to June 30, 2025.
Appendix D ACGME 6 Core Competencies
The Accreditation Council for Graduate Medical Education (ACGME) defines six core competencies as the foundational framework for assessing physician performance and professional development in graduate medical educationSwing (2007). These competencies—Patient Care, Medical Knowledge, Interpersonal and Communication Skills, Professionalism, Practice-Based Learning and Improvement, and Systems-Based Practice—capture complementary dimensions of clinical competence, ethical conduct, communication, lifelong learning, and system awareness, as summarized in Table 9. In this work, we adopt the ACGME core competencies as a competency-based lens to annotate and analyze ethical reasoning behaviors in LLM-generated clinical responses, enabling structured evaluation of model performance across clinically relevant professional dimensions.
Competency Abbrev. Summary (from Stanford GME / ACGME framing)
Patient Care PC Provide patient care that is compassionate, appropriate, and effective for treating health problems and promoting health.
Medical Knowledge MK Demonstrate knowledge of established and evolving biomedical, clinical, epidemiological, and social-behavioral sciences, and apply this knowledge to patient care.
Interpersonal and Communication Skills ICS Communicate and collaborate effectively with patients, families, and health professionals; includes cross-cultural communication, teamwork/leadership, consultative roles, and maintaining timely, legible records.
Professionalism P Commit to professional responsibilities and ethical principles; demonstrate integrity, respect, patient-first responsiveness, respect for privacy/autonomy, accountability, and sensitivity to diverse populations.
Practice-Based Learning and Improvement PBLI Investigate and evaluate one’s care, appraise and assimilate evidence, and continuously improve through self-evaluation and lifelong learning; includes QI methods, feedback incorporation, EBM skills, and use of IT for learning.
Systems-Based Practice SBP Be aware of and responsive to the larger health care system; effectively use system resources for optimal care; includes care coordination, cost awareness and risk-benefit analysis, advocacy for quality systems, interprofessional teamwork, and addressing system errors.
Table 9: Six ACGME Core Competencies summarized from the Stanford Graduate Medical Education (GME) “ACGME Core Competencies” page.
Appendix E Intraclass Correlation Coefficient (ICC) Calculation Formula
@@
To further illustrate scoring consistency, we additionally plotted scatter diagrams comparing the scores assigned to open-ended questions by three human experts and our Assessment Pipeline. The results are shown in Figure 6.
Figure 6: Scatter plots of open-ended question scores assigned by three human experts and our Assessment Pipeline.
Appendix F Annotation Interface
Figure 7 shows the annotation interface for MCQA-related tasks, while Figure 8 shows the interface for tasks related to open-ended question and rubrics.
Figure 7: Annoation inferface for MCQAs.
Figure 8: Annoation inferface for open-ended questions and rubrics.